How do you solve(square root of 3 + square root of 5)^2
Answer by stanbon(75887)

solve(square root of 3 + square root of 5)^2
------------
= (sqrt(3)+sqrt(5))^2
-----
= 3 + 2*sqrt(15) + 5
-----
= 7 + 2sqrt(15)
====================
Cheers,
Stan H.
============